OpenStudy (ilovecomputers1):

Express 3^4 = x as a logarithmic equation.

OpenStudy (mathstudent55):

\(\log_b x = y \iff b^y = x\) Start with the right side and end up with the left side.

OpenStudy (dtan5457):

For going between exponential form and log form, I use a made up formula called "BAE" 1.Base 2.Answer 3.Exponent \[3^4=x\] Base=3, so log 3 \[\log_{3} \] Answer=x, so log3x \[\log_{3}x \] Exponent=4 \[\log_3x=4\]

OpenStudy (dtan5457):

Vice versa. \[\log_{3}x=4 \] 1.Base=3 2. Answer=4 \[3^4\] 3. Exponent=x \[3^4=x\]

OpenStudy (ilovecomputers1):

@dtan5457 Hey sorry my computer froze up for a little bit, so I restarted it. That acronym helps a lot, thank you so much.

OpenStudy (dtan5457):

Log's are confusing. For me, that simple word of "Bae", is the only way I can remember it.